the length of a rectangle is 4 times the width. The perimeter is 35. find the width :
i triecthis formula but it isnt working out :
2( G x 4 ) + 2G =35

please help !

Answer by checkley77(12844)

L=4W
2L+2W=35
2(4W)+2W=35
8W+2W=35
10W=35
W=35/10
W=3.5 ANS. FOR THE WIDTH
L=4*3.5=14 ANS. FOR THE LENGTH.
PROOF:
2*14+2*3.5=35
28+7=35
35=35